# Geo4D: Leveraging Video Generators for Geometric 4D Scene Reconstruction

### v1

@@ 3.3 Multi-Modal Alignment
As noted, Geo4D predicts several non-independent geometric modalities. Furthermore, processing all frames of a long monocular video simultaneously with a video diffusion model is computationally prohibitive. Therefore, during inference, we use a temporal sliding window that segments the video into multiple overlapping video clips, with partial overlap to facilitate joining them. The goal of this section is to fuse the resulting multi-modal and multi-window data into a single, coherent reconstruction of the entire video.
@@
Given a video =$\mathcal{I}=\{\bm{I}^{i}\}_{i=1}^{N}$ with $N$ frames, we construct it into several video clips $\mathcal{G}=\{g^{k}\},k\in\mathcal{S}$, where each clip $g^{k}$ contains $V$ frames $\{\bm{I}^{i}\}_{i=k}^{k+V-1}$, and the set of starting indices is $\mathcal{S}=\{0,s,2s,\dots,\left\lfloor\frac{N-V}{s}\right\rfloor s\}\cup\{N-V\}$. Here, $s$ is the sliding window stride. The final term $\{N-V\}$ ensures that the last clip always includes the final frames of the video.
@@
First, given the predicted point maps $\bm{X}^{i,g}$ for each frame $i$ in each video clip $g\in\mathcal{G}$, we derive corresponding globally aligned point maps in world coordinates, as well as the relative camera motion and scale parameters. We denote these quantities with the $\mathrm{p}$ subscript to emphasize that they are inferred from the point map predictions. To do so, we extend the pair-wise global alignment loss from DUSt3R to a group-wise one:
@@
where $\lambda_{\text{p}}^{g}$ and $\bm{P}_{\text{p}}^{g}=[\mathbf{R}_{\text{p}}^{g}\mid\bm{\mathbf{\beta}}_{\text{p}}^{g}]$ denote the group-wise scale and transformation matrix that align the group-relative point maps $\bm{X}^{i,g}$ to the point maps $\bm{X}^{i}$ expressed in the global reference frame. We further parameterize each of these point maps as $\bm{X}_{uv}^{i}=\mathbf{R}_{\text{p}}^{i^{\top}}\mathbf{K}_{\text{p}}^{i^{-1}}{\bm{D}^{i}}^{-1}_{\text{p},uv}(u,v,1)+\bm{o}_{\text{p}}^{i}$ in terms of each camera’s calibration $\mathbf{K}^{i}_{\text{p}}$, world-to-camera rotation $\mathbf{R}^{i}_{\text{p}}$ and center $\bm{o}^{i}_{\text{p}}$ expressed in the global reference frame, and the disparity map $\bm{D}^{i}_{\text{p}}$. Substituting this expression into the loss function (4) and minimizing it, we can thus recover $\mathbf{K}^{i}_{\text{p}}$, $\mathbf{R}^{i}_{\text{p}}$, $\bm{o}^{i}_{\text{p}}$, $\bm{D}^{i}_{\text{p}}$, $\lambda_{\text{p}}^{g}$, $\bm{P}_{\text{p}}^{g}$ from the predicted point maps.
The steps above thus infer the disparity maps $\bm{D}_{\text{p}}^{i}$ from the point maps, but the model also predicts disparity maps $\bm{D}_{\text{d}}^{i}$ directly, where the d subscript denotes disparity prediction. We introduce the following loss to align them:
@@
$\mathcal{L}_{\text{c}}(\mathbf{R}_{\text{p}},\bm{o}_{\text{p}},\mathbf{R}_{\text{c}}^{g},\bm{\beta}_{\text{c}}^{g},\lambda_{\text{c}}^{g})=\sum_{g\in\mathcal{G}}\sum_{i\in g}(\left\|\mathbf{R}_{\text{p}}^{i^{\top}}\mathbf{R}^{g}_{\text{c}}\mathbf{R}^{i,g}_{\text{c}}-\bm{I}\right\|_{\mathrm{f}}$
$+\left\|\lambda_{\text{c}}^{g}\bm{o}_{\text{c}}^{i,g}+\bm{\beta}_{\text{c}}^{g}-\bm{o}_{\text{p}}^{i}\right\|_{2}),$ (8)
where $\bm{R}^{g}_{\text{c}},\bm{\beta}^{g}_{\text{c}},\lambda^{g}_{\text{c}}$ are learnable group-wise rotation matrix, translation vector and scale respectively to align the global camera trajectory $(\mathbf{R}_{p},\bm{o}_{p})$ and the predicted ones $(\mathbf{R}_{c},\bm{o}_{c})$. Following MonST3R [113], we also use a loss to smooth the camera trajectory:
@@
Abs Rel $\downarrow$ $\delta<1.25$uparrow Abs Rel $\downarrow$ $\delta<1.25$uparrow Abs Rel $\downarrow$ $\delta<1.25$uparrow
@@
Video depth & Camera pose Robust-CVD [32] 0.703 47.8 - - - -
@@
The model predicts point, disparity maps and ray map origins up to scale, as this cannot be uniquely determined from a monocular video. The disparity map is also recovered up to a translation, which discounts the focal length (this is sometimes difficult to estimate due to the dolly zoom effect). Likewise, the ray map origin is recovered up to shift, necessary to allow normalizing these maps.
@@ 4.1 Experimental Settings
Geo4D is trained exclusively on synthetic datasets, yet demonstrates strong generalization to real-world videos. Specifically, we use five synthetic datasets for training: Spring [50], BEDLAM [2], PointOdyssey [119], TarTanAir [93], and VirtualKitti [6]. See the Supp. Mat Tab. 5 for the details.
@@
Our Geo4D is initialized with the weight of Dynamicrafter [102] and trained using AdamW [48] with a learning rate of $1\times 10^{-5}$ and a batch size of 32. We use a progressive training strategy to improve convergence and stability, First, we train the model to generate a single geometric modality, i.e. the point maps, on a fixed resolution of $512\times 320$. Next, we introduce the multi-resolution training scheme to improve generalization and robustness, which includes various resolutions: $512\times 384$, $512\times 320$, $576\times 256$, $640\times 192$. Finally, we progressively add additional geometric modalities, i.e., the ray and depth maps. Training is conducted on 4 NVIDIA H100 GPUs with a total training time of approximately one week.
@@
As described in Sec. 3.2, given a $N$-frame video as input, we first split it into overlapping clips $\mathcal{G}$, each containing $V=16$ frames, with a stride of $s=4$. Each video clip, consisting of $V$ frames, is encoded and fed to the diffusion model to sample multi-modal 4D parameters $(\bm{X}^{i,g},\bm{D}^{i,g},\bm{r}^{i,g})$ for the video. For sampling, we use DDIM [77] with 5 steps. Finally, the alignment algorithm in Sec. 3.2 is used to fuse the clips into a globally coherent 4D reconstruction of the entire video.
@@ 4.2 Video Depth Estimation
Our hypothesis is that, despite being trained on synthetic data, our model can generalize well to out-of-distribution synthetic and real data as it is based on a pre-trained video diffusion model. To test this hypothesis, we evaluate our model on three benchmarks: Sintel [5] is a synthetic dataset that provides accurate depth annotations, covering diverse scenes with complex camera motion. KITTI [17] is a large driving dataset collected using stereo cameras and LiDAR sensors. Bonn [55] focuses on dynamic indoor scenes. To ensure fair comparisons, we follow the evaluation protocol used by MonST3R [113], where depth sequences are uniformly sampled from the datasets, extracting 50–110 frames per sequence for evaluation.
@@
We compare Geo4D to the state-of-the-art single-frame depth estimation (Marigold [28] and Depth-Anything-V2 [106]), video depth prediction (NVDS [95], ChronoDepth [70] and DepthCrafter [22]), and joint video depth and camera pose prediction (Robust-CVD [32], CasualSAM [116], and MonST3R [113]).
Figure 3: Qualitative results comparing Geo4D with MonST3R [113]. Attribute to our group-wise inference manner and prior geometry knowledge from pretrained video diffusion, our model successfully produces consistent 4D geometry under fast motion (first row) and deceptive reflection in the water (second row).
@@ Results.
As shown in Table 1, all versions of Geo4D outperform state-of-the-art methods by a large margin. This includes DepthCrafter [22] and MonST3R [113], the most recent video depth diffusion model and the dynamic extension of DUSt3d to dynamic scenes, respectively. Notably, while both Geo4D and DepthCrafter are based on the same video diffusion model (DynamiCrafter), our model outperforms DepthCrafter by Abs Rel by $24.0\%$ on Sintel and $17.3\%$ on KITTI datasets despite solving a more general problem. Qualitatively, Fig. 3 shows that Geo4D achieves more consistent results especially for fast moving objects.
@@ 4.3 Camera Pose Estimation
We evaluate the performance of our Geo4D on both synthetic Sintel [5] dataset and realistic TUM-dynamics [78] dataset. We follow the same evaluation protocol as in MonST3R [113]. Specifically, on Sintel, we select 14 dynamic sequences, and for TUM-dynamics, we sample the first 90 frames of each sequence with a temporal stride of 3. After aligning the predicted camera trajectory with the ground truth one with Umeyama algorithm, we calculate 3 commonly used metrics: Absolute Translation Error (ATE), Relative Translation Error (RPE-T), and Relative Rotation Error (RPE-R). We compare our method with other state-of-the-art discriminative methods, which jointly predict camera pose and depth, including Robust-CVD [32], CausalSAM [116], and MonST3R [113].
@@ Results.
To the best of our knowledge, Geo4D is the first method that uses a generative model to estimate the camera parameters in a dynamic scene. As shown in Tab. 2, compared to existing non-generative alternatives, we achieve much better camera rotation prediction (RPE-R) and comparable camera translation prediction (ATE and RPE-T).
@@
Point Map Depth Map Ray Map Point Map Depth Map Ray Map Abs Rel $\downarrow$ $\delta<1.25$uparrow ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
- - - - 0.232 71.3 0.335 0.076 0.731
- - 0.223 72.5 0.237 0.070 0.566
- - 0.211 73.4 - - -
- - - - 0.268 0.192 1.476
@@ 4.4 Qualitative Comparison
We compare Geo4D with the state-of-the-art MonST3R method on the DAVIS [23] dataset. Upgrading from pair-wise alignment as in MonsST3R to our group-wise alignment improves temporal consistency, leading to a more stable and globally coherent 4D reconstruction of point maps and camera trajectory, particularly in highly dynamic scenes. As shown in the top row in Fig. 3, Geo4D successfully tracks in 4D the racing car, whereas MonST3R struggles due to the rapid motion between pairs of images. Furthermore, probably because of the strong prior captured by the pre-trained video generative model, Geo4D correctly reconstructs the reflection of the flamingo in the water (second row in Fig. 3), whereas MonST3R misinterprets the reflection as a foreground object, resulting in incorrect depth.
@@
Abs Rel $\downarrow$ $\delta<1.25$uparrow ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
@@
Table 4: Ablation study for the temporal sliding window stride on the Sintel [5] dataset. Smaller strides are better.
@@
We compare Geo4D with state-of-the-art video depth predictors MonST3R [113] and DepthCrafter [22] on Sintel [5] dataset. Qualitatively, Geo4D produces more detailed geometry, for instance for the rope on the stick in the first row of Fig. 4, and a better spatial arrangement between different dynamic objects, as shown the second row of Fig. 4.
@@ 4.5 Ablation Study
We study the effect of multi-modality in Tab. 3. The three modalities, i.e. point map, depth map, and ray map, and can be used either at training or inference time, or both. The first two rows show that the diffusion model trained with point maps as a single modality performs worse in both video depth and camera pose estimation than the diffusion model trained with all three modalities. Therefore, the other two modalities, even if they can be seen as redundant, serve as additional supervisory signal during training, which improves the generalization ability of the diffusion model.
@@
We ablate the sliding window stride in Tab. 4. Results improve with a shorter stride, in part because this means that more windows and estimates are averaged, reducing the variance of the predictions by the denoising diffusion model, which is stochastic. We choose stride $s=4$ for our main results to balance the runtime and the performance. Note that MonST3R [113] requires 2.41 seconds to process one frame under the same setting. So, our method is 1.27 times faster than MonST3R [113].

### v2

@@ 3.3 Multi-Modal Alignment
As noted, Geo4D predicts several non-independent geometric modalities. Furthermore, processing all frames of a long monocular video simultaneously with a video diffusion model is computationally prohibitive. Therefore, during inference, we use a temporal sliding window that segments the video into multiple overlapping clips, with partial overlap to facilitate joining them. The goal of this section is to fuse the resulting multi-modal and multi-window data into a single, coherent reconstruction of the entire video.
@@
Given a video $\mathcal{I}=\{\bm{I}^{i}\}_{i=1}^{N}$ with $N$ frames, we divide it into several video clips $\mathcal{G}=\{g^{k}\}$, $k\in\mathcal{S}$, where each clip $g^{k}$ contains $V$ frames $\{\bm{I}^{i}\}_{i=k}^{k+V-1}$, and the set of starting indices is $\mathcal{S}=\{0,s,2s,\dots,\left\lfloor\frac{N-V}{s}\right\rfloor s\}\cup\{N-V\}$. Here, $s$ is the sliding window stride. The final term $\{N-V\}$ ensures that the last clip always includes the final frames of the video.
@@
First, given the predicted point maps $\bm{X}^{i,g}$ for each frame $i$ in each video clip $g\in\mathcal{G}$, we derive corresponding globally aligned point maps in world coordinates, as well as the relative camera motion and scale parameters. We denote these quantities with the $\mathrm{p}$ subscript to emphasize that they are inferred from the point map predictions. To do so, we extend the pairwise global alignment loss from DUSt3R to a group-wise one:
@@
where $\lambda_{\text{p}}^{g}$ and $\bm{P}_{\text{p}}^{g}=[\mathbf{R}_{\text{p}}^{g}\mid\bm{\mathbf{\beta}}_{\text{p}}^{g}]$ denote the group-wise scale and transformation matrix that align the group-relative point maps $\bm{X}^{i,g}$ to the point maps $\bm{X}^{i}$ expressed in the global reference frame. $\bm{\mathbf{\sigma}}_{uv}^{i,g}$ denotes the uncertainty of the point map for frame $i$ in group $g$ at pixel $(u,v)$. We further parameterize each of these point maps as $\bm{X}_{uv}^{i}=\mathbf{R}_{\text{p}}^{i^{\top}}\mathbf{K}_{\text{p}}^{i^{-1}}{\bm{D}^{i}}^{-1}_{\text{p},uv}(u,v,1)+\bm{o}_{\text{p}}^{i}$ in terms of each camera’s calibration $\mathbf{K}^{i}_{\text{p}}$, world-to-camera rotation $\mathbf{R}^{i}_{\text{p}}$, and center $\bm{o}^{i}_{\text{p}}$ expressed in the global reference frame, and the disparity map $\bm{D}^{i}_{\text{p}}$. Substituting this expression into the loss function (4) and minimizing it, we can thus recover $\mathbf{K}^{i}_{\text{p}}$, $\mathbf{R}^{i}_{\text{p}}$, $\bm{o}^{i}_{\text{p}}$, $\bm{D}^{i}_{\text{p}}$, $\lambda_{\text{p}}^{g}$, $\bm{P}_{\text{p}}^{g}$ from the predicted point maps.
The steps above infer the disparity maps $\bm{D}_{\text{p}}^{i}$ from the point maps, but the model also predicts disparity maps $\bm{D}_{\text{d}}^{i}$ directly, where the d subscript denotes disparity prediction. We introduce the following loss to align them:
@@
$\mathcal{L}_{\text{c}}(\mathbf{R}_{\text{p}},\bm{o}_{\text{p}},\mathbf{R}_{\text{c}}^{g},\bm{\beta}_{\text{c}}^{g},\lambda_{\text{c}}^{g})=\sum_{g\in\mathcal{G}}\sum_{i\in g}\Big{(}\left\|\mathbf{R}_{\text{p}}^{i^{\top}}\mathbf{R}^{g}_{\text{c}}\mathbf{R}^{i,g}_{\text{c}}-\bm{I}\right\|_{\mathrm{f}}$
$+\left\|\lambda_{\text{c}}^{g}\bm{o}_{\text{c}}^{i,g}+\bm{\beta}_{\text{c}}^{g}-\bm{o}_{\text{p}}^{i}\right\|_{2}\Big{)},$ (8)
where $\bm{R}^{g}_{\text{c}},\bm{\beta}^{g}_{\text{c}},\lambda^{g}_{\text{c}}$ are learnable group-wise rotation matrix, translation vector, and scale, respectively, to align the global camera trajectory $(\mathbf{R}_{p},\bm{o}_{p})$ and the predicted ones $(\mathbf{R}_{c},\bm{o}_{c})$. Following MonST3R [113], we also use a loss to smooth the camera trajectory:
@@
Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$
@@
Video depth & Camera pose Robust-CVD [32] 0.703 47.8
@@
The model predicts point maps, disparity maps, and ray map origins up to scale, as this cannot be uniquely determined from a monocular video. The disparity map is also recovered up to a translation, which discounts the focal length (this is sometimes difficult to estimate due to the dolly zoom effect). Likewise, the ray map origin is recovered up to a shift, necessary to allow normalizing these maps.
@@ 4.1 Experimental Settings
Geo4D is trained exclusively on synthetic datasets, yet demonstrates strong generalization to real-world videos. Specifically, we use five synthetic datasets for training: Spring [50], BEDLAM [2], PointOdyssey [119], TarTanAir [93], and VirtualKitti [6]. See the Supp. Mat Tab. 5 for details.
@@
Our Geo4D is initialized with the weights of DynamiCrafter [102] and trained using AdamW [48] with a learning rate of $1\times 10^{-5}$ and a batch size of 32. We use a progressive training strategy to improve convergence and stability. First, we train the model to generate a single geometric modality, i.e., the point maps, at a fixed resolution of $512\times 320$. Next, we introduce a multi-resolution training scheme to improve generalization and robustness, which includes various resolutions: $512\times 384$, $512\times 320$, $576\times 256$, $640\times 192$. Finally, we progressively add additional geometric modalities, i.e., the ray and depth maps. Training is conducted on 4 NVIDIA H100 GPUs with a total training time of approximately one week.
@@
As described in Sec. 3.2, given an $N$-frame video as input, we first split it into overlapping clips $\mathcal{G}$, each containing $V=16$ frames, with a stride of $s=4$. Each video clip is encoded and fed to the diffusion model to sample multi-modal 4D parameters $(\bm{X}^{i,g},\bm{D}^{i,g},\bm{r}^{i,g})$ for the video. For sampling, we use DDIM [77] with 5 steps. Finally, the alignment algorithm in Sec. 3.2 is used to fuse the clips into a globally coherent 4D reconstruction of the entire video.
@@ 4.2 Video Depth Estimation
Our hypothesis is that, despite being trained on synthetic data, our model can generalize well to out-of-distribution synthetic and real data, as it is based on a pre-trained video diffusion model. To test this hypothesis, we evaluate our model on three benchmarks: Sintel [5] is a synthetic dataset that provides accurate depth annotations, covering diverse scenes with complex camera motion. KITTI [17] is a large driving dataset collected using stereo cameras and LiDAR sensors. Bonn [55] focuses on dynamic indoor scenes. To ensure fair comparisons, we follow the evaluation protocol used by MonST3R [113], where depth sequences are uniformly sampled from the datasets, extracting 50–110 frames per sequence for evaluation.
@@
We compare Geo4D to state-of-the-art single-frame depth estimation methods (Marigold [28] and Depth-Anything-V2 [106]), video depth prediction (NVDS [95], ChronoDepth [70], and DepthCrafter [22]), and joint video depth and camera pose prediction (Robust-CVD [32], CausalSAM [116], and MonST3R [113]).
Figure 3: Qualitative results comparing Geo4D with MonST3R [113]. Attributed to our group-wise inference manner and prior geometry knowledge from pretrained video diffusion, our model successfully produces consistent 4D geometry under fast motion (first row) and deceptive reflection in the water (second row).
@@ Results.
As shown in Table 1, all versions of Geo4D outperform state-of-the-art methods by a large margin. This includes DepthCrafter [22] and MonST3R [113], the most recent video depth diffusion model and the dynamic extension of DUSt3R to dynamic scenes, respectively. Notably, while both Geo4D and DepthCrafter are based on the same video diffusion model (DynamiCrafter), our model outperforms DepthCrafter in Abs Rel by $24.0\%$ on Sintel and $17.3\%$ on KITTI, despite solving a more general problem. Qualitatively, Fig. 3 shows that Geo4D achieves more consistent results, especially for fast-moving objects.
@@ 4.3 Camera Pose Estimation
We evaluate the performance of Geo4D on both the synthetic Sintel [5] dataset and the realistic TUM-dynamics [78] dataset. We follow the same evaluation protocol as in MonST3R [113]. Specifically, on Sintel, we select 14 dynamic sequences, and for TUM-dynamics, we sample the first 90 frames of each sequence with a temporal stride of 3. After aligning the predicted camera trajectory with the ground truth using the Umeyama algorithm, we calculate three commonly used metrics: Absolute Translation Error (ATE), Relative Translation Error (RPE-T), and Relative Rotation Error (RPE-R). We compare our method with other state-of-the-art discriminative methods, which jointly predict camera pose and depth, including Robust-CVD [32], CausalSAM [116], and MonST3R [113].
@@ Results.
To the best of our knowledge, Geo4D is the first method that uses a generative model to estimate camera parameters in a dynamic scene. As shown in Tab. 2, compared to existing non-generative alternatives, we achieve much better camera rotation prediction (RPE-R) and comparable camera translation prediction (ATE and RPE-T).
@@
Point Map Disparity Map Ray Map Point Map Disparity Map Ray Map Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
0.232 71.3 0.335 0.076 0.731
0.223 72.5 0.237 0.070 0.566
0.211 73.4
0.268 0.192 1.476
@@ 4.4 Qualitative Comparison
We compare Geo4D with the state-of-the-art MonST3R method on the DAVIS [23] dataset. Upgrading from pairwise alignment as in MonST3R to our group-wise alignment improves temporal consistency, leading to a more stable and globally coherent 4D reconstruction of point maps and camera trajectory, particularly in highly dynamic scenes. As shown in the top row of Fig. 3, Geo4D successfully tracks the racing car in 4D, whereas MonST3R struggles due to the rapid motion between pairs of images. Furthermore, likely due to the strong prior captured by the pre-trained video generative model, Geo4D correctly reconstructs the reflection of the flamingo in the water (second row in Fig. 3), whereas MonST3R misinterprets the reflection as a foreground object, resulting in incorrect depth.
@@
Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
@@
Table 4: Ablation study for the temporal sliding window stride on the Sintel [5] dataset. There is a trade-off between performance and inference speed.
@@
We compare Geo4D with state-of-the-art video depth predictors MonST3R [113] and DepthCrafter [22] on the Sintel [5] dataset. Qualitatively, Geo4D produces more detailed geometry, for instance for the rope on the stick in the first row of Fig. 4, and a better spatial arrangement between different dynamic objects, as shown in the second row of Fig. 4.
@@ 4.5 Ablation Study
We study the effect of multi-modality in Tab. 3. The three modalities—point map, disparity map, and ray map—can be used either at training or inference time, or both. The first two rows show that the diffusion model trained with point maps as a single modality performs worse in both video depth and camera pose estimation than the diffusion model trained with all three modalities. Therefore, the other two modalities, even if they can be seen as redundant, serve as additional supervisory signals during training, which improves the generalization ability of the diffusion model.
@@
We ablate the sliding window stride in Tab. 4. Results improve with a shorter stride, in part because this means that more windows and estimates are averaged, reducing the variance of the predictions by the denoising diffusion model, which is stochastic. We choose stride $s=4$ for our main results to balance runtime and performance. Note that MonST3R [113] requires 2.41 seconds to process one frame under the same setting, so our method is 1.27 times faster than MonST3R [113].